OpenStudy (anonymous):

Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f + g)(5)?

OpenStudy (anonymous):

take \(f(5)=13\times 5^2\) and add \(g(5) = 7\times 5-5\)

OpenStudy (anonymous):

is the answer 355?